# Seeing Structural Failure Before it Happens: An Image-Based Physics-Informed Neural Network (PINN) for Spaghetti Bridge Load Prediction

### v1

@@ I Introduction
Physics Informed Neural Networks (PINNs) have emerged as a powerful class of models that integrate known physical laws—typically in the form of differential equations—into the training process of neural networks [1]. By embedding physics directly into the loss function, PINNs ensure that the learned solutions not only fit the data but also satisfy governing physical relationships. This hybrid approach has demonstrated superior generalization and robustness compared to purely data-driven methods, particularly in settings with limited or noisy data. PINNs have been successfully applied in a range of scientific and engineering domains, including fluid dynamics [2], biomedical engineering [3], geophysics [4], and heat transfer problems [5], where they have consistently outperformed conventional neural networks in producing physically consistent and interpretable results.
@@
In structural engineering, traditional analysis techniques—such as finite element methods—have long provided reliable results, but often at the cost of high computational expense and a reliance on detailed, sometimes hard-to-measure input parameters [6, 7]. As structural systems grow in complexity and scale, there is a growing need for predictive models that are not only computationally efficient but also capable of generalizing across varying configurations [8, 9]. While machine learning (ML) methods offer promising alternatives, their purely data-driven nature can lead to predictions that overlook essential physical constraints and may be physically implausible [10]. In this context, PINNs offer a compelling middle ground—blending data-driven flexibility with embedded domain knowledge to produce results that are both accurate and physically meaningful [11, 12].
@@
To explore the practical application of PINNs within the context of structural engineering, this study focuses on a simplified yet instructive case: spaghetti bridges. These small scale models serve as effective proxies for investigating the behavior of larger structural systems, which is why they are widely used in educational settings to demonstrate fundamental principles of structural mechanics [13]. Despite their simplicity, spaghetti bridges exhibit key mechanical behaviors found in full scale structures, including load transfer, stress distribution, and characteristic failure modes [14].
@@
In this work, we develop a physics-informed learning framework that predicts the weight of spaghetti bridges using structural parameters obtained either manually or through computer vision applied to bridge images. Beyond its educational utility—allowing students to evaluate their designs prior to physical testing—our framework demonstrates how integrating domain knowledge with machine learning can significantly improve prediction accuracy, especially in data-constrained settings. At the core of our approach is a novel hybrid neural architecture, the Physics Informed Kolmogorov Arnold Network (PIKAN), which merges universal function approximation theory with physical constraints derived from structural behavior. Despite working with a limited dataset of 15 physical bridges, augmented to 100 samples, our model delivers strong predictive performance, highlighting the potential of PINNs for lightweight structural modeling. We further develop a web-based interface for interactive parameter input and prediction, aiming to bridge the gap between theory and hands-on learning.

### v2

@@ I Introduction
Physics Informed Neural Networks (PINNs) have emerged as a powerful class of models that integrate known physical laws into the training process of neural networks [1]. By embedding physics directly into the loss function, PINNs ensure that the learned solutions not only fit the data but also satisfy governing physical relationships. This hybrid approach has demonstrated superior generalization and robustness compared to purely data-driven methods, particularly in settings with limited or noisy data. PINNs have been successfully applied in a range of scientific and engineering domains, including fluid dynamics [2], biomedical engineering [3], geophysics [4], and heat transfer problems [5], where they have consistently outperformed conventional neural networks in producing physically consistent and interpretable results.
@@
In structural engineering, traditional analysis techniques such as those based on finite element methods, have long provided reliable results, but often at the cost of high computational expense and a reliance on detailed, sometimes hard-to-measure input parameters [6, 7]. As structural systems grow in complexity and scale, there is a growing need for predictive models that are not only computationally efficient but also capable of generalizing across varying configurations [8, 9]. While machine learning (ML) methods offer promising alternatives, their purely data-driven nature can lead to predictions that overlook essential physical constraints and may be physically implausible [10]. In this context, PINNs offer a compelling middle ground that blends data-driven flexibility with embedded domain knowledge to produce results that are both accurate and physically meaningful [11, 12].
@@
To explore the practical application of PINNs within the context of structural engineering, we focus on spaghetti bridges which is a simplified yet instructive case. These small scale models serve as effective proxies for investigating the behavior of larger structural systems due to which they are widely used in educational settings to demonstrate fundamental principles of structural mechanics [13]. Despite their simplicity, spaghetti bridges exhibit key mechanical behaviors found in full scale structures, including load transfer, stress distribution, and characteristic failure modes [14].
@@
In this work, we develop a physics-informed learning framework that predicts the load of spaghetti bridges using structural parameters obtained either manually or through computer vision applied to bridge images. Beyond its educational utility that allows students to evaluate their designs prior to physical testing, our framework demonstrates how integrating domain knowledge with machine learning can significantly improve prediction accuracy particularly in data-constrained settings. At the core of our approach is a novel hybrid neural architecture, the Physics Informed Kolmogorov Arnold Network (PIKAN) that merges universal function approximation theory with physical constraints derived from structural behavior. Despite working with a limited dataset of 100 physical bridges, our model delivers strong predictive performance, highlighting the potential of PINNs for lightweight structural modeling. We further develop a web-based interface for interactive parameter input and prediction, aiming to bridge the gap between theory and hands-on learning.